If the difference between two complementary angles is 36 degrees, what are the measures of the two angles
Answer by checkley77(7172)

x+y=90
x-y=36 now addthem.
------------
2x=126
x=126/2
x=63 degrees ans.
63+y=90
y=90-63
y=27 ans.
Proof:
63-27=36
36=36